How many 6 fluid ounce servings are in 2 1/4 gallons of Apple Juice?
Answer #1

1 Gallon is = 128 fluid ounce

Number of 6 fluid ounces in 128 fluid ounces = 128/6 = 64/3

That is 1 gallon has 64/3 6 fluid ounce servings.

Therefore 2 1/4 = 9/4 gallons has how many 6 ounce fluid servings = (9/4) * (64/3) = 48 servings